# Towards Internet-Scale Training For Agents

### v1

@@ Abstract
The predominant approach for training web navigation agents gathers human demonstrations for a set of popular websites and hand-written tasks, but it is becoming clear that human data are an inefficient resource. We develop a pipeline to facilitate Internet-scale training for agents without laborious human annotations. In the first stage, an LLM generates tasks for 150k diverse websites. In the next stage, LLM agents complete tasks and produce trajectories. In the final stage, an LLM reviews the trajectories and judges their success. Language models are competitive with human annotators, detecting and filtering out harmful content with an accuracy of 97%, generating feasible tasks with an 89% rate, and judging successful trajectories with an 82.6% accuracy. Scaling the pipeline, agents based on Llama 3.1 70B solve 16.7% of tasks for 150k sites. Training on the data generated by our pipeline is competitive with training on human demonstrations. In data-limited settings derived from Mind2Web and WebLINX, we improve Step Accuracy by up to +89.5% and +122.1% respectively for agents trained on mixtures of data from our pipeline, and human data. When training agents with all available human data from these benchmarks, agents fail to generalize to diverse real sites, and adding our data improves their generalization by +149.0% for WebLINX and +156.3% for Mind2Web. Code will be available at: data-for-agents.github.io.
@@ 1 Introduction
The predominant approach to training LLM-based web navigation agents is to collect human demonstrations on a set of manually curated websites and tasks (Deng et al., 2023; Zhou et al., 2024b; Putta et al., 2024; Koh et al., 2024a; Liu et al., 2024; Lù et al., 2024; Rawles et al., 2023). Human data can be laborious to collect and becomes costly to scale as the breadth of skills that users require from language model agents grows. There are more than three-hundred million sites on the Internet (The Common Crawl Foundation, 2024), and the range of sites that researchers can manually prepare for human annotation represents a tiny fraction of the Internet. The key problem is that human data can become unreliable at scale. Human-written web navigation tasks are highly effective for popular sites, but reliability drops for sites with lower popularity due to annotators’ lack of familiarity. For these sites with lower popularity, which represent the majority of sites on the internet, human-written web navigation tasks are feasible only 40% of the time (discussed in Section 4), requiring a costly manual verification step. For the same pool of sites, language models improve feasibility rates to more than 80%. There is a growing need to automate data pipelines for the next generation of agents trained at internet scale. We address a key challenge by reducing the dependency on human data in the agent pipeline. We develop an automatic data pipeline that aims to facilitate Internet-Scale Training for Agents, which we shorten to InSTA, a pipeline that relies on synthetic web navigation tasks proposed, attempted, and evaluated by language models.
Figure 1: Overview of the proposed agent pipeline. We develop a pipeline for training web navigation agents at internet scale using tasks proposed, attempted, and evaluated by pretrained large language models. We generate 150k diverse tasks across 1M internet sites. Code for our data generation pipeline, and traces for agent rollouts will be available on our website: data-for-agents.github.io.
Our method operates in three stages. In the first stage, we employ a language model to propose candidate web navigation tasks for an agent to perform across 150k live sites on the Internet. Current works are limited to 200 popular sites (Lù et al., 2024; Rawles et al., 2023; Deng et al., 2023) that human annotators are likely to be familiar with. Language models help us scale to 1,000 times more sites than current efforts, with better coverage of real-world sites. One major consideration when scaling up training for agents is safety: building safe agents requires that we avoid sites with harmful, unsafe, or dangerous content. We evaluated the ability of language models to detect such content and aggressively filtered out 85% candidates from the initial 1M sites to 150k sites that are judged safe by language models. These models succeed at detecting safe content with an accuracy of 97%, compared to 75% human accuracy. With tasks generated across a safe and diverse set of websites, we proceed to run language model agents to attempt the generated tasks.
In the second stage of the pipeline, a language model agent attempts to complete tasks using a web browser. We provide the entire Playwright API to the agent, which operates the browser by generating function calls in the Playwright API, a standard choice in recent web navigation benchmarks (Chezelles et al., 2024; Du et al., 2024; Drouin et al., 2024). While existing LLM agents can be directly applied to our tasks, measuring their success presents another challenge: sites organize data in bespoke formats. Previous efforts circumvent this by enlisting human labelers to review attempts manually and judge their success (Lù et al., 2024; Deng et al., 2023), or by designing sites for benchmarking where formats are consistent (Zhou et al., 2024b; Koh et al., 2024a). However, neither approach scales efficiently, and human annotators may not be reliable for sites they are not familiar with in the tail of the distribution. In the third stage of the pipeline, we scale the evaluation using language models. We employ LLMs to judge (Lightman et al., 2024) whether a task is solved by the final timestep and obtain an accuracy of up to 93.1% in detecting successful trajectories for the most confident predictions. Llama-3.1-70B-Instruct solves 16.7% of zero-shot tasks with a judge confidence of conf = 1. In data-limited settings derived from Mind2Web and WebLINX, we improve Step Accuracy by up to +89.5% and +122.1%, respectively, when training agents on mixtures of data from our pipeline and human data. When training agents with all human data from these benchmarks, agents fail to generalize to diverse real sites, and adding our data improves their generalization by +149.0% for WebLINX and +156.3% for Mind2Web.
Our work develops InSTA, an automated data pipeline that aims to facilitate Internet-scale training for agents. Driven by pretrained language models, we generate web navigation tasks for 150k live websites, run LLM agents, and judge their success. By removing humans from the agent pipeline, we can scale and extend coverage towards sites where human data is less reliable. We present an analysis to improve safety and show that our data can outperform human demonstrations for training agents. Code for the pipeline will be available at: data-for-agents.github.io.
@@ 2 Related Works
There is an emerging paradigm in modern NLP using language models (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023a; b) as the backbone for agents (Andreas, 2022). These models show impressive reasoning capabilities (Bubeck et al., 2023; Zhong et al., 2024; Valmeekam et al., 2024) that allow them to generalize to downstream applications, such as web navigation, where text formats differ significantly from their training data. Search algorithms provide a secondary axis to improve the reasoning capabilities of language model agents (Yao et al., 2023b; Besta et al., 2024; Koh et al., 2024b; Zhou et al., 2024a) by providing an explicit algorithmic scaffold and allowing test-time compute to improve reasoning steps (Snell et al., 2024; Zhong et al., 2024). Although most of the work focuses on running language models as zero-shot agents, fine-tuning language models to improve their effectiveness as agents is becoming popular (Putta et al., 2024; Zeng et al., 2023; Zhang et al., 2023; Hong et al., 2023; Xie et al., 2024; Wang et al., 2024) as target benchmarks are becoming more difficult for zero-shot language models.
@@
There are a growing number of agent pipelines aimed at fine-tuning language models to improve their effectiveness as agents (Mitra et al., 2024; Zeng et al., 2023; Putta et al., 2024; Chen et al., 2023; Ou et al., 2024). However, driven by the limited data available, many such works train on data with significant overlap with their test environment, either with different tasks for the same environment configuration as the test setting (Deng et al., 2023), or the same tasks (Putta et al., 2024). We instead consider a setting where tasks and environment configurations are entirely separate between training and testing, creating a strong train-test split that follows recommended practice. This presents a challenge: the web navigation data for training LLM agents is limited (Deng et al., 2023; Lù et al., 2024). We address this challenge by increasing the scale of data collection and better coverage of the distribution of real-world sites. We train on diverse tasks generated by our pipeline and successfully transfer agents trained on our data to downstream benchmarks while maintaining a strong train-test split. Our training procedure resembles a modified FireAct (Chen et al., 2023), where language models jointly propose and evaluate tasks for agents.
@@
The majority of datasets for training web navigation agents rely on human annotators to create tasks (Zhou et al., 2024b; Koh et al., 2024a; Rawles et al., 2023), and provide demonstrations (Deng et al., 2023; Lù et al., 2024; Rawles et al., 2023; Shen et al., 2024). This approach has limits, as the breadth and diversity of tasks researchers can manually curate are dwarfed by the volume of sites on the Internet. There are more than three-hundred million sites on the Internet according to The Common Crawl Foundation (2024), and existing datasets are limited to about 150 popular sites that human annotators are already familiar with (Deng et al., 2023; Lù et al., 2024; Shen et al., 2024). There are hypothetical 1,000,000 times more data that could be available if we can efficiently harness this previously untapped resource. However, most sites are relatively obscure and human annotators are unreliable for sites they are not already familiar with. Finding suitable annotators becomes impractical at this massive scale, so we adapt language models to propose, attempt, and evaluate web navigation tasks. Although we are not the first to consider synthetic data for training agents (Gandhi et al., 2024; Ou et al., 2024; Setlur et al., 2024; Tajwar et al., 2024), we have developed a key approach to harness internet-scale data efficiently.
@@
Core to our pipeline is a language model evaluator. Using language models to judge the correctness of responses is becoming popular to improve accuracy for LLMs (Li et al., 2024), and applications include verifying reasoning steps (Zhang et al., 2024), rejection sampling (Snell et al., 2024; Sun et al., 2024), prioritizing frontier nodes in search algorithms (Zhou et al., 2024a; Koh et al., 2024b), filtering out harmful responses (Inan et al., 2023), providing feedback for response improvement (Madaan et al., 2023; Paul et al., 2024; Patel et al., 2024; Yuksekgonul et al., 2024), and providing ratings for alignment (Lee et al., 2024; Ouyang et al., 2024). Our use of language models to evaluate agent tasks is inspired by the generative verifier in Zhang et al. (2024), and modified from the multimodal verifier in He et al. (2024), where our language model predicts a confidence score that a task is solved, which is used to identify successful attempts.
@@ 3 Language Model Agents
Language model agents are a class of decision-making agents represented by $\pi_{\text{LLM}}(\mathbf{a}_{t}|\mathbf{s}_{t},\mathbf{c})$, a policy that processes multimodal observations $\mathbf{s}_{t}$ (from a web environment in our case) and predicts textual actions $\mathbf{a}_{t}$ to complete a task $\mathbf{c}$. Underneath this abstraction, a large language model (LLM) generates actions via the next-token prediction, conditioned on a system prompt $\mathbf{x}_{\text{sys}}$.
$\mathbf{a}_{t}=f^{\;\text{text}\to\text{act}}\big{(}\;\text{LLM}([\;\mathbf{x}_{\text{sys}},\mathbf{c},\text{Enc}(\mathbf{s}_{t})\;])\;\big{)}$ (1)
Environment representations for observations and actions typically differ from the expected input format of the language model (typically images and text), and functions are introduced that map the observations to a multimodal prompt $\text{Enc}(\cdot)$, and parse actions from the language model generated response $f^{\;\text{text}\to\text{act}}(\cdot)$. For web navigation, the environment state $\mathbf{s}_{t}$ is HTML DOM, and is often formatted as raw HTML code, an Accessibility Tree, Set-of-marks, or screenshots (Zhou et al., 2024b; Koh et al., 2024a; Chezelles et al., 2024; Shen et al., 2024). In this work, we employ an off-the-shelf HTML-to-Markdown parser to convert observations into a textual format. Action formats vary between works, and we build on Schick et al. (2023)’s function-calling framework, where a language model generates code that is parsed into a function name and corresponding arguments. Given a set of strings $L$, and a set of function argument values $G$, the set of actions $\mathcal{A}$ is:
$\mathcal{A}=\big{(}L_{\text{func}}\times(L_{\text{arg1}}\times G_{\text{arg1}})\times(L_{\text{arg2}}\times G_{\text{arg2}})\times\cdots\big{)}$ (2)
where $L_{\text{func}}$ is the set of function names in our API, and function arguments have a name and value ($L_{\text{arg1}}\times G_{\text{arg1}}$). We provide agent access to the entire Playwright API (Microsoft, 2024), a Microsoft-developed browser automation library that wraps around a Chromium web browser. The agent’s goal is to complete a web navigation task specified via a natural language instruction $\mathbf{c}\in L$, starting from an initial URL, and operating the browser via function calls to the Playwright API until the task is complete, after which point the agent calls stop with an optional answer:
@@
We prompt the language model backbone to generate actions wrapped in a JSON code block for straightforward parsing. The action parser $f^{\;\text{text}\to\text{act}}$ consists of a regex template that matches the first JSON code block, such as the example in Figure 1, followed by JSON decoding of the string contents within the code block. When parsing fails due to invalid syntax, we generate a new response until parsing succeeds. Equipped with a language model agent that makes calls to the Playwright API, we face a crucial roadblock that impedes scaling: obtaining large and diverse data.
@@ 4 Internet-Scale Task Generation
Building internet-scale agents requires a diverse scaffold of tasks and environment configurations beyond what can be attained via manually curated examples annotated by humans. We develop a pipeline to efficiently harness vast quantities of sites on the Internet that aims to facilitate Internet-Scale Training for Agents, shortened to InSTA. Our pipeline uses pretrained language models to generate, attempt, and evaluate synthetic web navigation tasks for a more diverse pool of sites than current efforts that rely on tasks designed by researchers (Deng et al., 2023; Zhou et al., 2024b; Putta et al., 2024; Koh et al., 2024a; Liu et al., 2024; Lù et al., 2024; Rawles et al., 2023; He et al., 2024). Human data is a valuable resource, but can be inefficient to gather. Our work shows that language models can be as accurate as human annotators. By removing human data from the agent pipeline, we can improve the safety and reliability of tasks, and efficiently scale task generation to 1M sites.

### v2

# InSTA: Towards Internet-Scale Training For Agents

@@ Abstract
The predominant approach for training web navigation agents is to gather human demonstrations for a set of popular websites and hand-written tasks, but it is becoming clear that human data is an inefficient resource. We develop a pipeline to facilitate internet-scale training for agents without laborious human annotations. In the first stage, an LLM annotates 150k sites with agentic tasks. In the next stage, LLM agents complete tasks and produce trajectories. In the final stage, an LLM filters trajectories by judging their success. Language models are powerful data curation tools, identifying harmful content with an accuracy of 97%, judging successful trajectories with an accuracy of 82.6%, and producing effective data. We train agents based on Qwen 3 1.7B that are competitive with frontier LLMs as web agents, while being smaller and faster. Our top agent reaches a success rate of 56.9%, outperforming the data collection policy Qwen 3 235B, a 235 times larger Llama 4 Maverick, and reaching 94.7% of the performance of Gemini 2.5 Flash. We are releasing code, models and data at: https://data-for-agents.github.io.
@@ 1 Introduction
The predominant approach for training LLM web navigation agents is to collect human demonstrations for a set of manually curated websites and tasks (Deng et al., 2023; Zhou et al., 2024b; Putta et al., 2024; Koh et al., 2024a; Liu et al., 2024; Lù et al., 2024; Rawles et al., 2023). Human data can be laborious to collect, and becomes costly to scale as the breadth of skills that users require from language model agents grows. There are more than 300M sites on the western internet according to The Common Crawl Foundation (2025), and the range of sites that researchers have annotated represents a tiny fraction of the vast available data. And crucially, the existing human data is static. There is a growing need to automate pipelines for training the next generation of language model agents in a dynamic internet-scale environment. This paper addresses the core challenge of building this environment—reducing dependence on human annotations. We develop an automatic pipeline that aims to facilitate internet-scale training for agents, which we refer to as the InSTA pipeline.
Figure 1: Overview of the InSTA pipeline. Our work unlocks a dynamic internet-scale environment that allows training small models to match frontier LLMs as agents, on a fraction of the budget. Starting from the top 1M sites on the internet, we annotate 150k sites with challenging agentic tasks, and release the entire pipeline, including code, models and an official huggingface dataset, on our website: https://data-for-agents.github.io.
The InSTA pipeline has three stages. In the first stage, we employ a language model task proposer that annotates 150k sites with live web navigation tasks for agents to perform. Existing works are limited to 200 popular websites (He et al., 2024; Lù et al., 2024; Rawles et al., 2023; Deng et al., 2023; Zhou et al., 2024c; Murty et al., 2025) that researchers have annotated manually, and a handful of synthetic websites (Zhou et al., 2024b; Koh et al., 2024a; Yao et al., 2023a). Our first goal in this paper is to improve coverage of real-world sites. To accomplish this, we cast a wide net. Starting from the top 1M sites on the internet ranked by popularity, our task proposer filters down to 150k sites that have safe content. Safety is critical when building autonomous agents, and our task proposer succeeds at detecting harmful content with an accuracy of 97%. Tasks are generated for sites marked as safe by the task proposer, and we run language model agents to complete the generated tasks. Agent progress is then fed back to the task proposer, which reviews trajectories and judge evaluations in order to assign a harder task based on the latest content of the website, closing the loop.
Scaling the task generation loop, we annotate 150k diverse sites with challenging agentic tasks, and release an official huggingface dataset: data-for-agents/insta-150k-v2. Motivated by the importance of internet data to progress in modern deep learning, our second goal in this paper is an internet-scale data flywheel for training LLM agents. We approach this by harnessing LLMs as data curation tools. After generating tasks, the pipeline employs pretrained LLMs as agents to complete tasks and produce trajectories, which are filtered by a judge to select the best data. Agents control a virtual web browser and produce reasoning traces that contain function calls to interact with and navigate live webpages. The judge produces a reasoning trace that considers whether a trajectory is successful, and scores the agent on a continuous scale from 0 to 1. We scale the pipeline to create a large reasoning dataset for multimodal agents, with 2.2M screenshots, 2.2M traces for actions, and 150k traces for the judge.
Our data unlocks great potential in small language models as agents. We train a series of models based on Qwen 3 1.7B on varying scales of data from the InSTA pipeline, and match the performance of frontier LLM agents, on a fraction of the budget. Our top agent has a success rate of 56.9%, outperforming the data collection policy Qwen 3 235B, beating a 235 times larger Llama 4 Maverick, and reaching 94.7% of the performance of Gemini 2.5 Flash. To share our progress, we are releasing the entire pipeline, including code, models and data, on our website: https://data-for-agents.github.io.
@@ 2 Related Works
There is an emerging paradigm in modern NLP using language models (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023a, b) as the backbone for agents (Andreas, 2022). These models show impressive reasoning capabilities (Bubeck et al., 2023; Zhong et al., 2024; Valmeekam et al., 2024) that allow them to generalize to downstream applications, such as web navigation, where observations differ from LLM training data. Search algorithms provide a secondary axis to improve the reasoning capabilities of language model agents (Yao et al., 2023b; Besta et al., 2024; Koh et al., 2024b; Zhou et al., 2024a) by providing an explicit algorithmic scaffold and allowing test-time compute to improve reasoning steps (Snell et al., 2024; Zhong et al., 2024). Although most recent work focuses on running language models as zero-shot agents, fine-tuning language models to improve their effectiveness as agents is becoming popular (Putta et al., 2024; Zeng et al., 2023; Zhang et al., 2023; Hong et al., 2023; Xie et al., 2024; Wang et al., 2024; Zhou et al., 2024c; Murty et al., 2025) as target benchmarks are becoming more difficult to solve zero-shot.
@@
There are a growing number of agent pipelines aimed at fine-tuning language models to improve their effectiveness as agents (Mitra et al., 2024; Zeng et al., 2023; Putta et al., 2024; Chen et al., 2023; Ou et al., 2024; Zhou et al., 2024c; Murty et al., 2025). However, driven by the limited data available, many such works train on data with significant overlap with their test environment, either with different tasks for the same environment configuration as the test setting (Deng et al., 2023; Zhou et al., 2024c; Murty et al., 2025), or even the same tasks (Putta et al., 2024). We consider a setting where tasks and environment configurations (i.e. websites) are entirely separate between training and testing, creating a strong train-test split. This presents a challenge: human data for training LLM agents is limited (Deng et al., 2023; Lù et al., 2024). We address this challenge by reducing dependence on human data in agent pipelines. We employ LLMs as data curation tools to automatically design challenging tasks, and select the best training data. Our pipeline allows us to train small models that match frontier LLMs on Web Voyager (He et al., 2024), without using any data from Web Voyager. Contrast this with methods that train primarily on data from Web Voyager (Zhou et al., 2024c; Murty et al., 2025), and may not transfer to other benchmarks (Xue et al., 2025).
@@
Datasets for training web navigation agents typically rely on human annotators to create tasks (Zhou et al., 2024b; Koh et al., 2024a; Rawles et al., 2023), and provide demonstrations for tasks (Deng et al., 2023; Lù et al., 2024; Rawles et al., 2023; Shen et al., 2024). However, the amount of data researchers have annotated represents a tiny fraction of the available internet data. There are more than 300M sites on the internet according to The Common Crawl Foundation (2025), yet existing datasets are limited to 200 popular sites that human annotators are familiar with (Deng et al., 2023; Lù et al., 2024; Shen et al., 2024). Human data can be laborious to collect, and becomes costly to scale as the capabilities users require from agents grows. And crucially, human data is static. Our work moves away from fixed datasets for training agents, and towards a dynamic internet-scale environment that grows with an ever-changing internet. We are not the first to build an environment (Zhou et al., 2024b; Koh et al., 2024a; Yao et al., 2023a; He et al., 2024), nor are we the first to consider synthetic data (Gandhi et al., 2024; Ou et al., 2024; Setlur et al., 2024; Tajwar et al., 2024), but we have solved a key challenge that unlocks the internet as the largest environment for agents.
@@
Using LLMs to judge the correctness of responses is becoming popular to refine LLM predictions (Li et al., 2024), including to verify reasoning steps (Zhang et al., 2024), rejection sample (Snell et al., 2024; Sun et al., 2024), prioritize frontier nodes in search algorithms (Zhou et al., 2024a; Koh et al., 2024b), filter out harmful responses (Inan et al., 2023), provide feedback for response improvement (Madaan et al., 2023; Paul et al., 2024; Patel et al., 2024; Yuksekgonul et al., 2024), and provide ratings for alignment (Lee et al., 2024; Ouyang et al., 2024). Our use of language models to evaluate agents is inspired by Generative Verifiers (Zhang et al., 2024), and the multimodal verifier in He et al. (2024). One difference is our verifier predicts a reasoning trace that scores the agent from 0 to 1, which helps us rank trajectories to select the best data.
@@ 3 Language Model Agents
Language model agents are a class of decision-making agents represented by $\pi_{\text{LLM}}(\mathbf{a}_{t}|\mathbf{s}_{t},\mathbf{c})$, a policy that processes multimodal observations $\mathbf{s}_{t}$ (from a virtual web browser in our case) and predicts textual actions $\mathbf{a}_{t}$ to complete a task $\mathbf{c}$. Underneath this abstraction, a large language model (LLM) generates actions via the next-token prediction, conditioned on a system prompt $\mathbf{x}_{\text{agent}}$.
$\mathbf{a}_{t}=f^{\;\text{text}\to\text{act}}(\;\text{LLM}(\;\mathbf{x}_{\text{agent}},\mathbf{c},\text{Enc}(\mathbf{s}_{t})\;)\;)$ (1)
Environment representations for observations and actions typically differ from the expected input format of the language model (typically images and text), and functions are introduced that map the observations to a multimodal prompt $\text{Enc}(\cdot)$, and parse actions from the language model generated response $f^{\;\text{text}\to\text{act}}(\cdot)$. For web navigation, the environment state $\mathbf{s}_{t}$ is HTML DOM, and is often formatted as raw HTML code, an Accessibility Tree, Set-of-marks, or screenshots (Zhou et al., 2024b; Koh et al., 2024a; Chezelles et al., 2024; Shen et al., 2024). We built a fast Markdown parser that converts webpage observations into a compact readable format (refer to the code). Action formats vary between works, and we build on Schick et al. (2023)’s function-calling framework, where a language model generates code that is parsed into a function name and corresponding arguments. Given sets of function and argument names $L_{i}$, and sets of argument values $G_{i}$, the action space $\mathcal{A}$ is:
$\mathcal{A}=L_{\text{func}}\times(L_{\text{arg1}}\times G_{\text{arg1}})\times(L_{\text{arg2}}\times G_{\text{arg2}})\times\cdots\times(L_{\text{argN}}\times G_{\text{argN}})$ (2)
where $L_{\text{func}}$ is the set of function names on the page object in the Playwright API (Microsoft, 2024), and function arguments have a name and value ($L_{\text{arg1}}\times G_{\text{arg1}}$) corresponding to the Playwright API. We allow the agent access to call arbitrary functions in Playwright (Microsoft, 2024), a Microsoft-developed browser automation library that wraps a headless web browser. The agent’s goal is to complete a web navigation task specified via a natural language instruction $\mathbf{c}\in L$, starting from an initial URL, and operating the browser via function calls to the Playwright API until the desired task is complete, after which the agent calls the stop function and provides a final response:
@@
We prompt the agent to produce a reasoning trace of a desired length (ablated in Figure 10) that contains function call actions as JSON in a fenced code block. To parse actions from the response, we employ a regex template that matches the first JSON code block, and a JSON decoder $f^{\;\text{text}\to\text{act}}(\cdot)$ to parse the contents within the code block. When parsing fails due to invalid syntax, we allow the agent to generate a second response. Equipped with a language model agent that makes function calls with the Playwright API, we may consider the crucial task of obtaining large and diverse data.
@@ 4 Internet-Scale Task Generation
Training the next generation of LLM agents requires a large and diverse set of websites and tasks beyond what researchers have annotated so far (Deng et al., 2023; Zhou et al., 2024b; Koh et al., 2024a; Liu et al., 2024; Lù et al., 2024; Rawles et al., 2023; He et al., 2024). We develop an approach to efficiently annotate vast numbers of sites from diverse sections of the internet with agentic tasks. Our approach introduces two important desiderata: (1) it should not rely on human annotations, and (2) tasks should derive from a feedback process that deeply explores the environment.